OpenStudy (anonymous):

A 364-mL container holds 0.146 g of Ne and an unknown amount of Ar at 35°C and a total pressure of 624 mmHg. Calculate the moles of Ar present.

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

Ok, so in this container, we have a mixture of neon and argon. However, both gases experience the same volume and temperature. We will need to find the total amount of moles in the mixture.

OpenStudy (ipwnbunnies):

Ideal Gas Equation: PV = nRT, rewrite for n, which is # of total moles of mixture. n = (PV)/(RT) So, in this problem, we'll have to use a different form of R. R = 0.08206 L-atm/mol-K

OpenStudy (anonymous):

ok i know v= .364 T= 308.15K P=.8211 and now R=0.08206

OpenStudy (ipwnbunnies):

We have to do some conversions with the given quantities. V = 364 mL (1L/1000mL) = 364 * 10^-3 L P = 624 mmHg (1atm/760 mmHg) = ?? atm [Please calculate this lol] T = 35 degrees C + 273.15= ?? K [calc this too]

OpenStudy (anonymous):

way ahead of ya lol. but idk what to do with the .146 grams cause thats Ne and we need to find Ar

OpenStudy (ipwnbunnies):

Did you find the total # of moles? I'll let you do all the calculations.

OpenStudy (ipwnbunnies):

Hold on, I just wanna know the total moles first.

OpenStudy (ipwnbunnies):

I'll just skip ahead then, assuming we have total moles... We need to find how many moles of Ne are in the mixture. moles = (mass)/(molar mass). Can you do that?

OpenStudy (anonymous):

i did .146/20.179 and it is .0072 moles

OpenStudy (anonymous):

now what?

OpenStudy (ipwnbunnies):

Great! So, moles of Ar = total moles - moles of Ne! That is all. :)

OpenStudy (anonymous):

so um im confused. what do i do now?

OpenStudy (ipwnbunnies):

Well, did you find the total moles from the beginning? n = (PV)/(RT). Plug in those values that you typed out.

OpenStudy (ipwnbunnies):

n = how many moles of gas are in the container.

OpenStudy (anonymous):

n=84.61 but why did i do .146/20.179?

OpenStudy (ipwnbunnies):

That's how many moles of neon are in the container. It's part of the total amount of gas in the container.

OpenStudy (ipwnbunnies):

And the total number of moles is actually 0.118 moles. Soooooo moles of Argon in the container = total moles (n) - moles of Neon

OpenStudy (ipwnbunnies):

So, mol Ar = 0.118 mol - 0.0072 mol = 0.111 mol Ar

OpenStudy (anonymous):

waiiiiit whered you get 0.118?

OpenStudy (ipwnbunnies):

0.118 was the total number of moles in the mixture. You found the wrong number (you said it was 84.62 mol I believe).

OpenStudy (anonymous):

._. really? my hard work down the drain lol

OpenStudy (anonymous):

(.08206)(308.15) -------------- = n (0.8211)(0.364) isn't this right?

OpenStudy (ipwnbunnies):

Lol, just go back to the top: n = [(0.8211 atm)(0.364 L)]/[(0.08206 L-atm/mol-K)(308.15 K)]

OpenStudy (ipwnbunnies):

Nooo, you flipped the fraction. n = (PV)/(RT)

OpenStudy (anonymous):

heheh oooops! ok but .111 isnt the answer right?

OpenStudy (ipwnbunnies):

Yes, 0.118 is the total amount of moles of gas in the container. We found the amount of moles of neon in the container, 0.0072 mols. The amount of Ar is the difference between them: 0.118 mol - 0.0072 mol = 0.111 mol of Argon

OpenStudy (anonymous):

it says i got it wrong lol

OpenStudy (ipwnbunnies):

Did you type out the right values at the very beginning?

OpenStudy (anonymous):

.0046 is the answer

OpenStudy (ipwnbunnies):

Hold on.

OpenStudy (anonymous):

ya .0118-.0072=.0046 :D

OpenStudy (ipwnbunnies):

Ffs, I just got that too. I'm so sorry. :/

OpenStudy (ipwnbunnies):

Can't believe I got 0.118 twice before. Unreal.

OpenStudy (anonymous):

lol dont be! i corrected it! see you taught me and i caught a mistake xD THANK YOU SOOOO MUCH! YOURE A LIFESAVER

OpenStudy (ipwnbunnies):

No prob lol.